# Can the Waymo Open Motion Dataset Support Realistic Behavioral Modeling? A Validation Study with Naturalistic Trajectories

### v1

@@ Abstract
The Waymo Open Motion Dataset (WOMD) has become a popular resource for data-driven modeling of autonomous vehicles (AVs) behavior. However, its validity for behavioral analysis remains uncertain due to proprietary post-processing, the absence of error quantification, and the segmentation of trajectories into 20-second clips. This study examines whether WOMD accurately captures the dynamics and interactions observed in real-world AV operations. Leveraging an independently collected naturalistic dataset from Level 4 AV operations in Phoenix, Arizona (PHX), we perform comparative analyses across three representative urban driving scenarios: discharging at signalized intersections, car-following, and lane-changing behaviors. For the discharging analysis, headways are manually extracted from aerial video to ensure negligible measurement error. For the car-following and lane-changing cases, we apply the Simulation–Extrapolation (SIMEX) method to account for empirically estimated error in the PHX data and use Dynamic Time Warping (DTW) distances to quantify behavioral differences. Results across all scenarios consistently show that behavior in PHX falls outside the behavioral envelope of WOMD. Notably, WOMD underrepresents short headways and abrupt decelerations. These findings suggest that behavioral models calibrated solely on WOMD may systematically underestimate the variability, risk, and complexity of naturalistic driving. Caution is therefore warranted when using WOMD for behavior modeling without proper validation against independently collected data.
@@ 1 Introduction
Autonomous vehicles (AVs), once considered futuristic, are now operating commercially in many cities. Over the past 40 years, numerous studies have proposed optimized AV car-following [1, 2] and lane-changing [3, 4] models aimed at improving traffic flow efficiency. The majority of these studies focus on lower levels of autonomy (e.g., Adaptive Cruise Control - ACC - and automated lane-changing). Moreover, most of these studies rely on simplifying assumptions about the AV software stack. For example, detailed vehicle dynamics models are rarely incorporated when analyzing string stability, and perception challenges are often reduced to structured noise models. Such simplifications are in stark contrast to the realities of AV deployment, as the industry moves towards the deployment of AVs at high levels of automation.
@@
In practice, AV development has been shaped more by safety, reliability, regulatory compliance, and user comfort than by theoretical optimality (often at the cost of efficiency). Traditional robotic-based architectures used in commercial AVs leave little room for any simplifications, as they must ensure robust operation under diverse and uncertain conditions, even if that requires suboptimal decisions due to computational constraints and safety requirements. Compounding this divergence, recent foundation model-based approaches largely prioritize reproducing observed human driving behavior over optimizing for traffic efficiency, abandoning the notion of optimality in favor of compatibility with human driver expectations.
@@
Accordingly, several recent studies have deviated from focusing on optimality and utilized real-world data from AV operations to model AV behavior [5], assess its impacts on traffic flow dynamics [6], and characterize the impacts of AVs on human behavior in their vicinity [7, 8]. The majority of these studies have utilized one of three available datasets from real-world AV operations at high levels of automation: Waymo Open Motion Dataset (WOMD) [9, 10], Lyft level-5 Open Dataset [11], and Argoverse Dataset [12]. Considering that the Lyft Level-5 Open Dataset and Argoverse Dataset are no longer being updated, WOMD has become a dominant dataset for such analyses. Although a wide range of topics have been covered in studies that utilized WOMD, they all have one common underlying assumption: WOMD is accurate and suitable for behavioral and traffic flow research. Unfortunately, this dataset offers no avenue for independent accuracy validation. Focusing on the WOMD, there exist two sources of error in this dataset: (1) Ego vehicle poses are derived from IMU/GNSS and are therefore an order of magnitude more precise than the LiDAR, camera, and radar fusion used for surrounding objects. Researchers cannot quantify the error structure because the raw sensor data and the technical details for data processing are withheld, and (2) all published trajectories are the outcome of a proprietary off-board tracking pipeline that implicitly applies temporal smoothing across several past and future frames for regressing the current bounding box. While this filtering is advantageous for motion planning benchmarks, as it delivers ”clean” inputs with reduced measurement noise, it can potentially suppress exactly the discontinuous elements that characterize real human driving, for example, stop–and-go oscillations in queuing, start-up reaction at signalized intersections, and the lateral jerk during lane-changing behaviors.
Thus, it is essential to know the extent of inaccuracies in the WOMD as any model calibrated based on these smoothed trajectories will understate variability in acceleration, gap acceptance, and reaction latency, ultimately biasing conclusions about safety and efficiency. Note that the aforementioned potential issues are in addition to the challenges associated with 20-second-long trajectories (WOMD partitions continuous traffic into 20-second snippets). This design maximizes scenario diversity for prediction challenges (for which this dataset is prepared) but lacks the temporal context needed for the calibration of microscopic models. Critical variables such as lead–lag history and multi-lane interactions fall outside the retained time window, making it impossible to accurately estimate car following or lane choice models.
@@
Accordingly, this study aims to answer one fundamental question regarding WOMD: Can this dataset support behavioral and traffic flow research? To answer this question, the observed behavior in the WOMD is compared to an independently collected naturalistic dataset from Level 4 AV operations in Phoenix, AZ (PHX). The analysis focus on three representative driving scenarios: intersection discharging, car-following, and lane-changing. For the intersection discharging case, headways are manually extracted from high-resolution aerial videos in the PHX dataset with negligible error. These manually extracted headways are then compared against those extracted from WOMD trajectories at intersections, which rely on vehicle pose estimates and map alignment. For the car-following and lane-changing scenarios, we analyze trajectories from both datasets with error-aware statistical tools to assess their alignment. In particular, we employ the Simulation–Extrapolation (SIMEX) procedure [13] to remove the effect of empirically estimated measurement noise in the PHX dataset. This correction allows for a fair comparison of behavior patterns such as decelerating-to-stop responses and lane-change interactions. By quantifying cross-dataset differences and comparing them against the internal variability of WOMD, we assess whether the naturalistic behaviors observed in the PHX dataset fall within the behaviors captured in WOMD.
@@ 2.1.1 Data collection
The first step is to perform a comprehensive data collection. The goal was to capture uninterrupted AV driving in real-world conditions. The two main locations selected for data collection are Phoenix and Chandler, AZ. The selected locations, specifically downtown Phoenix, allow for the existence of a series of challenging real-world scenarios that the AVs have to navigate. Data collection was performed in late January 2023, in clear, sunny weather conditions, between 11 am and 5 pm. Sunny weather conditions usually mean increased shadows in the collected data, which makes data post-processing steps more challenging. The data collection method used in this work is based on the Moving Aerial Videography [14, 15], which includes a helicopter mounted with an 8K resolution camera that collects birds-eye view footage at 500 feet altitude by following the vehicles of interest throughout the study duration. High-quality, information-rich interactions are clipped into smaller segments for later processing and analysis. The behavior of AVs is captured during straight road segment movements, left/right turns, at intersections, and within several other driving scenarios. The AVs present in this data collection are Level-4 fully driverless vehicles that include absolutely no human intervention throughout the study. They are capable of car-following, lane-changing, navigating intersections, and stopping for pedestrians and other objects they may face. Several runs of data collection are performed, where a complete helicopter run consists of approximately 12 miles of road navigated on the ground.
@@ 2.1.2 Trajectory extraction
Vehicle trajectory information is later extracted from the recorded videos using the methodology explained in [14, 15]. The method begins by extracting high-resolution raw images from the video recordings, which are recorded at 30 frames per second. Next, a few representative frames are selected to create a reference image of the study area, which is necessary to transform the coordinates of each image into a global frame of reference. The next steps include object detection and object tracking. Subsequently, image stabilization is performed to ultimately transform each frame to the global coordinate system.Once transformed into a global frame, trajectories are extracted.
@@ 2.1.3 Error Estimation
To generate ground truth travel distances, we manually identified 40 distinct control points along a 2.2-km southbound corridor (Figure 1). These points were selected using clear visual anchors on the roadway, such as the tip of turn arrows, crosswalk stripes, and other unambiguous road markings. The control points were georeferenced using satellite imagery with 0.3-meter resolution and embedded UTM coordinates [16]. The absolute differences in UTM Northing and Easting coordinates between control points define the longitudinal and lateral distances traveled, respectively.
For each vehicle trajectory, we reviewed raw video frames to determine the exact time at which the vehicle center passed each control point. Frame numbers were converted into timestamps based on the video’s sampling rate (30 fps), and all valid time-point pairs were identified. To obtain a representative distribution of travel distances, we employed stratified random sampling without replacement across four trajectory duration quartiles: intervals below 16.9 seconds, between 16.9–44.8 seconds, 44.8–80.6 seconds, and above 80.6 seconds. A total of 60 pairs were selected. For each pair, the measured longitudinal and lateral distances were compared to ground truth to obtain the 2D distance-traveled error vector.
The resulting distribution of travel distance error was modeled using a bivariate normal distribution. As shown in figure 2, the final error distribution, based on 50 travel distance samples, yielded the following parameters: mean longitudinal error $\mu_{x}=0.276$ m, mean lateral error $\mu_{y}=0.006$ m, standard deviations $\sigma_{x}=1.075$ m and $\sigma_{y}=0.530$ m, and a moderate negative correlation $\rho_{xy}=-0.291$. Mardia’s multivariate skewness and kurtosis tests returned $p=0.104$ and $p=0.580$, respectively, indicating that the null hypothesis of joint normality cannot be rejected. From this travel distance error model, we derived the distribution of spacing error between two vehicles by taking the difference of two independent errors. Assuming independence, the spacing error follows a distribution $g\sim\mathcal{N}(0,\;2\sigma_{x}^{2})=\mathcal{N}(0,\;2.309\;\text{m}^{2})$.
@@ 2.2 Waymo Open Motion Dataset
The motion dataset offers 20-second segments sampled at 10 Hz, representing a variety of scenarios along with their corresponding map data. The dataset includes details such as a unique scenario ID, unique tracking IDs for each object, object types (i.e., vehicles, pedestrians, and cyclists), whether a vehicle is Waymo or not, and attributes like object position (x, y, z), dimensions (length, width, height), heading, and velocity. For each scenario, the dataset also offers detailed map information, including lane centers, lane boundaries, and road boundaries, represented as 3D polylines or polygons. Further details about the motion dataset is available at (https://waymo.com/open/data/motion/).

### v2

@@ Abstract
The Waymo Open Motion Dataset (WOMD) has become a popular resource for data-driven modeling of autonomous vehicles (AVs) behavior. However, its validity for behavioral analysis remains uncertain due to proprietary post-processing, the absence of error quantification, and the segmentation of trajectories into 20-second clips. This study examines whether WOMD accurately captures the dynamics and interactions observed in real-world AV operations. To achieve this objective, this paper introduces a robust, error-aware validation methodology for assessing whether a public vehicle-trajectory dataset is suitable for driving behavior and traffic-flow research using an independently collected ground-truth reference dataset with a characterized error distribution. Accordingly, leveraging an independently collected naturalistic dataset from Level 4 AV operations in Phoenix, Arizona (PHX), we perform comparative analyses across three representative urban driving scenarios: discharging at signalized intersections, car-following, and lane-changing behaviors. For the discharging analysis, headways are manually extracted from aerial video to ensure negligible measurement error. For the car-following and lane-changing cases, we apply the Simulation–Extrapolation (SIMEX) method to account for empirically estimated error in the PHX data and use Dynamic Time Warping (DTW) distances to quantify behavioral differences. Results across all scenarios consistently show that behavior in PHX falls outside the behavioral envelope of WOMD. Notably, WOMD underrepresents short headways and abrupt decelerations. These findings suggest that behavioral models calibrated solely on WOMD may systematically underestimate the variability, risk, and complexity of naturalistic driving. Caution is therefore warranted when using WOMD for behavior modeling without proper validation against independently collected data.
@@ 1 Introduction
Autonomous vehicles (AVs), once considered futuristic, are now operating commercially in many cities. Over the past 40 years, numerous studies have proposed optimized AV car-following (Talebpour and Mahmassani, 2016; Mahdinia et al., 2021) and lane-changing (An et al., 2023; An and Talebpour, 2023) models aimed at improving traffic flow efficiency. The majority of these studies focus on lower levels of autonomy, including Adaptive Cruise Control (ACC) and automated lane-changing. Moreover, most of these studies rely on simplifying assumptions about the AV software stack. For example, detailed vehicle dynamics models are rarely incorporated when analyzing string stability, and perception challenges are often reduced to structured noise models. Such simplifications are in stark contrast to the realities of AV deployment, as the industry moves towards the deployment of AVs at high levels of automation.
@@
In practice, AV development has been shaped more by safety, reliability, regulatory compliance, and user comfort than by theoretical optimality (often at the cost of efficiency). Traditional robotic-based architectures used in commercial AVs leave little room for any simplifications, as they must ensure robust operation under diverse and uncertain conditions, even if that requires suboptimal decisions due to computational constraints and safety requirements. Compounding this divergence, recent foundation model-based approaches largely prioritize reproducing observed human driving behavior over optimizing for traffic efficiency, abandoning the notion of optimality in favor of compatibility with human driver expectations. Examples include large Transformer-based motion forecasting models such as Wayformer (Nayakanti et al., 2022), vectorized end-to-end driving stacks such as VAD (Jiang et al., 2023; Chen et al., 2024b), and emerging surveys of foundation models for autonomous driving perception and planning (Sathyam and Li, 2025; Gao et al., 2024). These systems are trained on vast datasets and optimize multi-task objectives, but they inherit any measurement and post-processing biases present in the underlying datasets.
@@
Accordingly, several recent studies have deviated from focusing on optimality and utilized real-world data from AV operations to model AV behavior (Hu et al., 2022), assess its impacts on traffic flow dynamics (Hu et al., 2023a), and characterize the impacts of AVs on human behavior in their vicinity (Zhang and Talebpour, 2024; Rahmati et al., 2019). The majority of these studies have utilized one of three available datasets from real-world AV operations at high levels of automation: Waymo Open Motion Dataset (WOMD) (Ettinger et al., 2021; Chen et al., 2024a), Lyft level-5 Open Dataset (Houston et al., 2021), and Argoverse Dataset (Chang et al., 2019). Considering that the Lyft Level-5 Open Dataset and Argoverse Dataset are no longer being updated, WOMD has become a dominant dataset for such analyses. Although a wide range of topics have been covered in studies that utilized WOMD, they all have one common underlying assumption: WOMD is accurate and suitable for behavioral and traffic flow research. Unfortunately, this dataset offers no avenue for independent accuracy validation.
Despite its breadth and popularity, the Waymo Open Motion Dataset (WOMD) has several structural limitations when used for microscopic behavioral analysis. First, WOMD is constructed from 20 s segments for interactions and then further cropped to 9 s scenarios for the Interaction Prediction Challenge ***The details for Waymo Open Dataset Interaction Prediction Challenge can be found at https://waymo.com/open/challenges/2025/interaction-prediction/, rather than preserving long, continuous runs of naturalistic driving. This design is ideal for motion-forecasting challenges, but it truncates car-following and lane-changing histories and removes information about how drivers adapt over multiple signal cycles, merges, and shockwaves. As a result, key state variables such as long-range headway history, multi-episode stopping and re-acceleration, or sequence of lane changes cannot be reconstructed, which limits the calibration and validation of dynamic behavioral models that inherently depend on a longer temporal context.
Second, the released WOMD scenarios are effectively outdated. Since the initial public release in 2021, the dataset has contained the same 103,354 20 s scenarios; subsequent versions only add new modalities, including enhanced maps, LiDAR point clouds, camera embeddings, and traffic signals for these existing clips, without introducing new driving episodes. Consequently, the trajectories predominantly reflect autonomous-driving policies deployed before 2021, whereas the commercial operations of the major Level 4 ride-hailing provider in Phoenix have undergone substantial software and policy updates since then. In other words, even if WOMD trajectories were perfectly measured, they still represent an outdated slice of autonomous-vehicle behavior.
Third, all trajectories in WOMD are the output of an offboard perception pipeline that explicitly aggregates information over a point-cloud sequence, including both past and future frames, to estimate the current bounding box of an object and generate temporally smooth 3D tracks (Qi et al., 2021; Ettinger et al., 2021). This offline smoothing greatly reduces perception noise and tracking fragmentation but also suppresses exactly the kinds of discontinuities that are critical to behavioral and traffic-flow modeling, such as start-up delays at signals, stop-and-go oscillations in queues, abrupt hazard-induced braking, or lateral oscillation during lane changes. Any model calibrated directly on these smoothed trajectories will tend to understate variability in acceleration, gap acceptance, and reaction times. Moreover, recent work has shown that even non-trajectory portions of WOMD, such as traffic signal states, exhibit substantial missing and inaccurate entries with 71.7% of signal states unknown or missing in the raw data and requiring substantial post hoc imputation and correction (Yan et al., 2026).
@@
Accordingly, this study aims to answer one fundamental question regarding WOMD: Can this dataset support behavioral and traffic flow research? To answer this question, the observed behavior in the WOMD is compared to an independently collected naturalistic dataset from Level 4 AV operations in Phoenix, AZ (PHX). The analysis focuses on three representative driving scenarios: intersection discharging, car-following, and lane-changing. For the intersection discharging case, headways are manually extracted from high-resolution aerial videos in the PHX dataset with negligible error. These manually extracted headways are then compared against those extracted from WOMD trajectories at intersections, which rely on vehicle pose estimates and map alignment. For the car-following and lane-changing scenarios, we analyze trajectories from both datasets with error-aware statistical tools to assess their alignment. In particular, we employ the Simulation–Extrapolation (SIMEX) procedure (Cook and Stefanski, 1994) to remove the effect of empirically estimated measurement noise in the PHX dataset. This correction allows for a fair comparison of behavior patterns such as decelerating-to-stop responses and lane-change interactions. By quantifying cross-dataset differences and comparing them against the internal variability of WOMD, we assess whether the naturalistic behaviors observed in the PHX dataset fall within the behaviors captured in WOMD.
It is important to note that this is not a perfectly controlled experiment. The PHX trajectories are collected in 2023 from a specific corridor in Phoenix under normal driving scenarios, whereas WOMD aggregates self-driving operations from multiple U.S. cities before 2021. Moreover, commercial AV behavior evolves rapidly. Even if WOMD trajectories were free of measurement and post-processing error, a nontrivial portion of the discrepancy we observe may reflect (i) changes in the underlying driving policy over time, (ii) differences in roadway geometry and traffic composition across cities, and (iii) the use of different sensing and smoothing pipelines, for example, aerial videography versus on-board, proprietary perception and tracking. Accordingly, our goal is not to attribute discrepancies solely to a single source, but rather to quantify whether recent, naturalistic Level 4 operation in Phoenix lies within the behavioral envelope spanned by WOMD, and in doing so, to critically assess the validity of using WOMD directly for driving behavior studies.
It is also important to note that several studies have examined the accuracy and internal consistency of vehicle trajectory datasets, particularly in the context of the Next Generation Simulation (NGSIM) program (Punzo et al., 2011; Montanino and Punzo, 2013, 2015). Among the most influential contributions, Punzo et al. (2011) proposed a comprehensive framework for assessing trajectory data quality based on kinematic consistency diagnostics, including jerk distributions, acceleration smoothness, and spectral properties of vehicle motion. In that work, a Kalman-filtered reconstruction of the original trajectories was used as a reference to quantify inconsistency in the raw NGSIM data. While this approach established important benchmarks for identifying measurement noise and post-processing artifacts, it relies on internally reconstructed trajectories derived from the same data source rather than an external ground-truth reference. As a result, the assessment primarily evaluates internal consistency and physical plausibility, rather than absolute validity with respect to independently observed vehicle motion. Subsequent studies adopting similar principles have likewise focused on smoothing-based corrections and internal diagnostics (Hu et al., 2022, 2023b), reflecting the limited availability of independent, high-resolution ground truth in large-scale trajectory datasets. In contrast, the present study advances trajectory dataset validation by leveraging an independently collected naturalistic dataset with empirically characterized measurement error, enabling an external and error-aware assessment of a public autonomous-vehicle trajectory dataset.
Considering the abovementioned discussions, this study makes two contributions: First, we introduce a robust methodology to assess the validity of a vehicle trajectory dataset. Second, leveraging a self-collected naturalistic robotaxi dataset (PHX) with empirically characterized error, to our knowledge, we provide the first independent validation of WOMD using third-party data (serving as the ground truth) rather than relying on WOMD’s own post-processed outputs as a reference.
@@ 2.1.1 Data collection
The first step is to perform a comprehensive data collection. The goal was to capture uninterrupted AV driving in real-world conditions. The two main locations selected for data collection are Phoenix and Chandler, AZ. The selected locations, specifically downtown Phoenix, allow for the existence of a series of challenging real-world scenarios that the AVs have to navigate. Data collection was performed in late January 2023, in clear, sunny weather conditions, between 11 am and 5 pm. Sunny weather conditions usually mean increased shadows in the collected data, which makes data post-processing steps more challenging. The data collection method used in this work is based on the Moving Aerial Videography (Ammourah et al., 2024; Talebpour et al., 2024), which includes a helicopter mounted with an 8K resolution camera that collects birds-eye view footage at 500 feet altitude by following the vehicles of interest throughout the study duration. High-quality, information-rich interactions are clipped into smaller segments for later processing and analysis. The behavior of AVs is captured during straight road segment movements, left/right turns, at intersections, and within several other driving scenarios. The ego vehicles targeted in this data collection are driverless robotaxis offered as a public ride service by a single major provider of automated mobility in Phoenix and Chandler.According to public statements and SAE J3016, these services operate without an on-board safety driver within a geo-fenced operational design domain that corresponds to Level 4 automated driving capability. Throughout the study runs, these vehicles operated without any human intervention inside the vehicle cabin. They performed everyday driving tasks such as car-following, lane-changing, navigating signalized intersections, and yielding to pedestrians and other road users. All surrounding traffic in the PHX dataset consists of conventional human-driven vehicles sharing the roadway with the ego robotaxis.Several runs of data collection are performed, where a complete helicopter run consists of approximately 12 miles of road navigated on the ground.
@@ 2.1.2 Trajectory extraction
Vehicle trajectory information is later extracted from the recorded videos using the methodology explained in (Ammourah et al., 2024; Talebpour et al., 2024). The method begins by extracting high-resolution raw images from the video recordings, which are recorded at 30 frames per second. Next, a few representative frames are selected to create a reference image of the study area, which is necessary to transform the coordinates of each image into a global frame of reference. The next steps include object detection and tracking, where vehicles in each frame are detected using a RetinaNet-based detector that was retrained based on the data collected in this study (Hosseini et al., 2022). Object tracking includes assigning a unique ID to each vehicle using data association and track maintenance (Hosseini et al., 2022). Subsequently, image stabilization is performed to ultimately transform each frame to the global coordinate system.Once transformed into a global frame, trajectories are extracted.
@@ 2.1.3 Error Estimation
To generate ground truth travel distances, we manually identified 40 distinct control points along a 2.2-km southbound corridor (Figure 1). The $(x,y)$ coordinates in the PHX dataset are defined in a global frame constructed by transforming the aerial frames into a high-resolution reference image, which is geometrically matched to a geo-referenced satellite image of the study area. This matching aligns the axes of the reference image with UTM Easting and Northing (Snyder, 1987), so that motion in $x$ and $y$ closely matches the change in UTM coordinates. The downtown Phoenix experiment area is laid out as an almost perfectly rectilinear grid with very small curvature within the data collection area, so roadway centerlines are approximately parallel or orthogonal to these axes. For each corridor, we therefore define the longitudinal direction along the predominant axis of travel and the lateral direction along the orthogonal axis.
For each vehicle trajectory, we reviewed raw video frames to determine the exact time at which the vehicle center passed each control point. Frame numbers were converted into timestamps based on the video’s frame rate at 30 fps, and all valid time-point pairs were identified. The resulting ground truth is therefore accurate up to the limits imposed by the video frame rate and pixel resolution in the 8K imagery on the order of 0.0333 seconds and 0.3 meters. This construction makes the PHX dataset a trajectory set with empirically characterized error: the residuals between aerially extracted trajectories and the manually labeled reference distances are precisely what our error model captures. To obtain a representative distribution of travel distances, we employed stratified random sampling without replacement across four trajectory duration quartiles: intervals below 16.9 seconds, between 16.9–44.8 seconds, 44.8–80.6 seconds, and above 80.6 seconds. This stratification reflects the expectation that tracking error may increase with the length and duration of the interval; it also prevents very short hops from dominating the error sample. A total of 60 pairs were selected. For each pair, the measured longitudinal and lateral distances were compared to ground truth to obtain the 2D distance-traveled error vector.
The resulting distribution of travel distance error was modeled using a bivariate normal distribution. As shown in figure 2, the final error distribution, based on 60 travel distance samples, yielded the following parameters: mean longitudinal error $\mu_{x}=0.276$ m, mean lateral error $\mu_{y}=0.006$ m, standard deviations $\sigma_{x}=1.075$ m and $\sigma_{y}=0.530$ m, and a moderate negative correlation $\rho_{xy}=-0.291$. Mardia’s multivariate skewness and kurtosis tests returned $p=0.104$ and $p=0.580$, respectively, indicating that the null hypothesis of joint normality cannot be rejected. From this travel distance error model, we derived the distribution of spacing error between two vehicles by taking the difference of two independent errors. Assuming independence, the spacing error follows a distribution $g\sim\mathcal{N}(0,\;2\sigma_{x}^{2})=\mathcal{N}(0,\;2.309\;\text{m}^{2})$.
@@ 2.2 Waymo Open Motion Dataset
The motion dataset offers 20-second segments sampled at 10 Hz, representing a variety of scenarios along with their corresponding map data. The dataset includes details such as a unique scenario ID, unique tracking IDs for each object, object types (i.e., vehicles, pedestrians, and cyclists), whether a vehicle is Waymo or not, and attributes like object position (x, y, z), dimensions (length, width, height), heading, and velocity. For each scenario, the dataset also offers detailed map information, including lane centers, lane boundaries, and road boundaries, represented as 3D polylines or polygons. Further details about the motion dataset are available at (https://waymo.com/open/data/motion/).
A key distinction is that the PHX trajectories are derived from a stabilized, top-down aerial viewpoint, whereas WOMD trajectories are produced by an ego-centric perception and tracking stack mounted on the self-driving vehicle and by offboard processing. As a result, PHX errors are dominated by object-tracking uncertainty in orthorectified imagery, whereas WOMD errors are driven by on-board sensor noise, occlusions, and proprietary temporal smoothing.
In summary, WOMD provides a broad but temporally and spatially heterogeneous snapshot of AV operations, while the PHX dataset offers a smaller but more homogeneous sample of everyday robotaxi driving in a specific urban corridor in early 2023. The combination of these two views allows us to assess whether typical maneuvers observed in up-to-date commercial service are well represented in the public dataset.